surrounding a living cell consists of an inner and an outer wall that are separated by a small space. Assume that the membrane acts like a parallel plate capacitor in which the effectiv of 6.9 x 10-6 C/m2, (a) what is the magnitode of the electric field within the cel membrane? (b) Find the magnitude of the electric force that would be exerted on a potassium ion (K+; charge +e) placed inside the membrane.
30 GO exerted on the object by the electric feld
Answer #1

Problem 36 :

The electric field at any point due to an infinite sheet of charge density sigma is given by

  E= 2E0

Where,  epsilon _{0} = 8.85 imes 10^{-12} , , C^{2}/Nm^{2}

And the electric field inside the capacitor made by infinite sheets is given by

E = rac{sigma}{ epsilon _{0}}

Given,

  σ = 6.9 × 10-0 C/㎡

(a) So, magnitude of electric field within the membrane is

6.9 × 10-6 8.85 × 10-12 = 0.78 × 106 N/C

(b) The force on a charge q inside an electric field E is given by

F= qE

As the potassium ion has charge ' +e ' , so the magnitude of force exerted on the potassium due to the field of the membrane is

13 F = 1.602 × 10-19 × 0.78 × 10°-1.25 × 10-

Problem 30 :

(a) Given, the object has charge of +10.0 x 10-6 C and the force exerted on it is 50.0 x 10-6 N .

We have,

F= qE

So, 50.0 × 10-o-10.0 × 10-0 × E

or,   E = 5 , , N/C

And the direction of electric field will be same as the direction of force i.e. due east.

(b) Given, the object has charge of -25.0 x 10-6 C and the force exerted on it is 10.0 x 10-6 N .

So, F 10.0 × 10-0 -25.0 x 10-6 g -0.4 УС

The negative sign is indicating the the direction of electric field is in the opposite direction of the force.

Hence, the electric field has magnitude 0.4 N and has a direction towards east (opposite of west) .

For any doubt please comment and please give an up vote. Thank you.